If sec(α)<0 and the sin(α)=15/17

What is the exact value of cot(α)
1
Expert's answer
2020-11-17T06:37:18-0500

"sec(\alpha)=\frac 1 {cos(\alpha)}" . if "sec(\alpha)<0" then "cos(\alpha)<0" .

"cos(\alpha)=-\sqrt{1-sin^{2}(\alpha)}=-\sqrt{1-(\frac {15} {19})^{2}}=-\sqrt{\frac{64} {289}}=-\frac 8 {17}"

"cot(\alpha)=\frac {cos(\alpha)} {sin(\alpha)}=-\frac 8 {15}"